OpenStudy (anonymous):

@satellite73 f(x)= x/1-

OpenStudy (anonymous):

x=y/1-3y correct?

OpenStudy (anonymous):

that is the first step, if you are looking for the inverse then you have to solve for \(y\)

OpenStudy (anonymous):

i keep getting 4x x-3xy=y then divide the one side by y it cancels and you are left with 2x

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

you have to put all the \(y\) on one side and everything else on the other

OpenStudy (anonymous):

\[x-3xy=y\] lets put all the \(y\) on the right \[x=y+3xy\]

OpenStudy (anonymous):

factor out the \(y\) and write \[x=y(1+3x)\]

OpenStudy (anonymous):

then divide to get \(y\) by itself \[\frac{x}{1+3x}=y\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

these take a little practice, but they are not that hard once you do a couple of them

OpenStudy (anonymous):

@satellite73 ive been able to continue my homework on the same type of problems thanks a lot